4. Buckeye Company has provided the following data for maintenance cost:

Prior Year Current Year
  Machine hours 14,100        17,000       
  Maintenance cost $28,200        $31,970       

Maintenance cost is a mixed cost with variable and fixed components. The fixed and variable components of maintenance cost are closest to:
Answer #1

Variable cost per machine hour = (Highest activity cost - Lowest activity cost)/(Highest activity - Lowest activity)

= (31,970-28,200)/(17,000-14,100)

= 3,770/2,900

= $1.3 per machine hour

Fixed cost = Highest activity cost - Highest activity x Variable cost per machine hour

= 31,970 - (17,000 x 1.3)

= 31,970-22,100

= $9,870

Kindly give a positive rating if you are satisfied with this solution and please ask if you have any query.

Thanks